OpenStudy (taepoohwae):

Can someone please simplify this with steps? (a² + 4a²) + (a³ - 5a² + 7a)

OpenStudy (anonymous):

only terms with the same exponent can be combined so x + x = 2x using this knowledge a^2 + 4a^2 = 5a^2 moving on to the next group none have the same exponent so they are fine, but you still have the 5a^2 from the first group of parentheses 5a^2+ a^3 - 5a^2+7a 5a^2 - 5a^2 = 0 so final answer is a^3 + 7a

OpenStudy (anonymous):

*also when adding those with the same exponent you add the coefficients. so 1x + 1x = 2x because 1 + 1 = 2 as with the a^2 terms 1a^2 + 4a^2 = 5a^2 because 1 + 4 = 5

OpenStudy (taepoohwae):

Thanks a bunch!